OpenStudy (anonymous):

if a is 40% of b, then b exceeds a by what percent of a?

OpenStudy (anonymous):

please give explanations too , please

OpenStudy (lgbasallote):

a is 40% of b so.. \[a = \frac{40}{100} \times b\]

OpenStudy (anonymous):

yes i got that far

OpenStudy (lgbasallote):

wait..im thinking this through too :D

OpenStudy (lgbasallote):

so \[b = \frac{100}{40}\times a\]

OpenStudy (anonymous):

whyd you flip it? :0

OpenStudy (lgbasallote):

i multiplied both sides by 100/40 to get rid of 40/100 on the right side does that make sense?

OpenStudy (anonymous):

OHH i see it

OpenStudy (lgbasallote):

\[b = \frac{40 + 60}{40} \times a\] do you see this too?

OpenStudy (anonymous):

hmm can you explain

OpenStudy (lgbasallote):

i just broke 100 into 40 + 60

OpenStudy (anonymous):

oh haha

OpenStudy (lgbasallote):

because of this \[b = (\frac{40}{40} + \frac{60}{40} ) \times a\]

OpenStudy (lgbasallote):

still seeing it?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

\[b = (1 + \frac 32) a\] what about this?

OpenStudy (anonymous):

yup

OpenStudy (lgbasallote):

im pretty sure you see this too \[b = a+ \frac 32 a\]

OpenStudy (anonymous):

kkkk

OpenStudy (lgbasallote):

next would be \[b = a + \frac{3\times 50}{2 \times 50} a\] \[b = a + \frac{150}{100} a\] so b exceeds a by 150% of a i guess..

OpenStudy (anonymous):

thanks!

OpenStudy (lgbasallote):

welcome ^_^